OpenStudy (anonymous):

How do you write 71/25 as a percentage?

OpenStudy (anonymous):

You multiply it by 100. O.O \[71/25 * 100 = 7100/25 = 284 percent\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

@tanvi: percentages can be calculated out of 100 basis , so it can't be >100

OpenStudy (anonymous):

It can. 2/1 = 200%. If it's more than the denominator, then it's MORE than 100% of the original value.

OpenStudy (anonymous):

yeah , the calculation you showed is correct, but express it as %

OpenStudy (anonymous):

I know, but the 'equation" thing doesn't show it :P

OpenStudy (anonymous):

lol